# OpenXAI: Towards a Transparent Evaluation of Model Explanations

### v1

@@ E.5 Additional Results on LR models
[…]
Table 11: Stability of explanation methods on the Heloc dataset with LR model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS
@@
[…]
Table 12: Stability of explanation methods on the Adult Income dataset with LR model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS
@@
[…]
E.6 Results on ANN models
Table 13: Predicted faithfulness results on the SynthGauss dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
[…]
Table 14: Predicted faithfulness results on the German dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
[…]
Table 15: Predicted faithfulness results on the Heloc dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
[…]
Table 16: Predicted faithfulness results on the Adult Income dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
[…]
Figure 6: Fairness analysis of PGU metric on the German Credit dataset with ANN model. Shown are average and standard error values for majority (male) and minority (female) subgroups.
Figure 7: Fairness analysis of PGU metric on the Adult Income dataset with ANN model. Shown are average and standard error values for majority (male) and minority (female) subgroups.
Table 17: Stability of explanation methods on the SynthGauss dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@ E.6 Results on ANN models
[…]
Table 18: Stability of explanation methods on the German Credit dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@
[…]
Table 19: Stability of explanation methods on the Heloc dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@
[…]
Table 20: Stability of explanation methods on the Adult Income dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@
[…]
Table 21: Stability of explanation methods on the COMPAS dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@
[…]
Table 22: Stability of explanation methods on the GMSC Credit dataset with ANN model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS ROS
@@
[…]
-$3±\pm0$.03
[…]
$3±\pm0$.03
[…]

### v5

@@ D.6 Remaining Results on LR models
0.503$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.877$\pm$0.00
1.000$\pm$0.00
0.854$\pm$0.00
0.950$\pm$0.00
0.008$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.895$\pm$0.00
1.000$\pm$0.00
0.868$\pm$0.00
0.973$\pm$0.00
0.155$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.555$\pm$0.01
1.000$\pm$0.00
0.544$\pm$0.00
0.822$\pm$0.00
0.052$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.210$\pm$0.01
1.000$\pm$0.00
0.181$\pm$0.01
0.433$\pm$0.01
0.078$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.555$\pm$0.01
1.000$\pm$0.00
0.544$\pm$0.00
0.822$\pm$0.00
0.027$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.210$\pm$0.01
1.000$\pm$0.00
0.181$\pm$0.01
0.433$\pm$0.01
0.122$\pm$0.00
0.090$\pm$0.00
0.090$\pm$0.00
0.100$\pm$0.00
0.090$\pm$0.00
0.100$\pm$0.00
0.090$\pm$0.00
0.047$\pm$0.00
0.096$\pm$0.00
0.096$\pm$0.00
0.086$\pm$0.00
0.096$\pm$0.00
0.086$\pm$0.00
0.096$\pm$0.00
Table 17: Ground-truth and predictive faithfulness results on the German Credit dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
0.500$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.465$\pm$0.00
1.000$\pm$0.00
0.505$\pm$0.00
0.977$\pm$0.00
-0.000$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.122$\pm$0.01
1.000$\pm$0.00
-0.001$\pm$0.01
0.995$\pm$0.00
0.129$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.171$\pm$0.01
1.000$\pm$0.00
0.172$\pm$0.01
0.971$\pm$0.00
0.014$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.026$\pm$0.00
1.000$\pm$0.00
0.025$\pm$0.00
0.837$\pm$0.01
0.067$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.171$\pm$0.01
1.000$\pm$0.00
0.171$\pm$0.01
0.971$\pm$0.00
0.007$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.026$\pm$0.00
1.000$\pm$0.00
0.025$\pm$0.00
0.837$\pm$0.01
0.049$\pm$0.00
0.035$\pm$0.00
0.035$\pm$0.00
0.040$\pm$0.00
0.035$\pm$0.00
0.040$\pm$0.00
0.035$\pm$0.00
0.011$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
Table 18: Ground-truth and predictive faithfulness results on the COMPAS dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
0.504$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.622$\pm$0.00
1.000$\pm$0.00
0.614$\pm$0.00
0.977$\pm$0.00
0.009$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.320$\pm$0.01
1.000$\pm$0.00
0.287$\pm$0.01
0.981$\pm$0.00
0.209$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.311$\pm$0.01
1.000$\pm$0.00
0.311$\pm$0.01
0.999$\pm$0.00
0.137$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.142$\pm$0.01
1.000$\pm$0.00
0.134$\pm$0.01
0.997$\pm$0.00
0.107$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.311$\pm$0.01
1.000$\pm$0.00
0.310$\pm$0.01
0.999$\pm$0.00
0.067$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.142$\pm$0.01
1.000$\pm$0.00
0.134$\pm$0.01
0.997$\pm$0.00
0.072$\pm$0.00
0.055$\pm$0.00
0.055$\pm$0.00
0.070$\pm$0.00
0.055$\pm$0.00
0.069$\pm$0.00
0.055$\pm$0.00
0.029$\pm$0.00
0.058$\pm$0.00
0.058$\pm$0.00
0.033$\pm$0.00
0.058$\pm$0.00
0.033$\pm$0.00
0.058$\pm$0.00
Table 19: Ground-truth and predictive faithfulness results on the GMSC dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
0.499$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.499$\pm$0.00
1.000$\pm$0.00
0.495$\pm$0.00
0.909$\pm$0.00
-0.006$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
-0.006$\pm$0.01
1.000$\pm$0.00
-0.053$\pm$0.01
0.920$\pm$0.00
0.197$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.124$\pm$0.01
1.000$\pm$0.00
0.140$\pm$0.01
0.954$\pm$0.00
0.098$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.067$\pm$0.01
1.000$\pm$0.00
0.085$\pm$0.01
0.953$\pm$0.00
0.098$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.124$\pm$0.01
1.000$\pm$0.00
0.140$\pm$0.01
0.954$\pm$0.00
0.049$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.067$\pm$0.01
1.000$\pm$0.00
0.085$\pm$0.01
0.953$\pm$0.00
0.026$\pm$0.00
0.012$\pm$0.00
0.012$\pm$0.00
0.023$\pm$0.00
0.012$\pm$0.00
0.023$\pm$0.00
0.012$\pm$0.00
0.009$\pm$0.00
0.024$\pm$0.00
0.024$\pm$0.00
0.013$\pm$0.00
0.024$\pm$0.00
0.013$\pm$0.00
0.024$\pm$0.00
Table 20: Stability results on the HELOC dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
11.74$\pm$0.00
2.75$\pm$0.02
1.25$\pm$0.01
2.75$\pm$0.02
10.00$\pm$0.03
9.68$\pm$0.02
9.86$\pm$0.03
14.10$\pm$0.02
1.89$\pm$0.13
-0.42$\pm$0.02
3.64$\pm$0.05
12.27$\pm$0.04
12.01$\pm$0.04
12.82$\pm$0.04
Table 21: Stability results on the Adult Income dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
12.95$\pm$0.00
3.75$\pm$0.02
2.61$\pm$0.01
3.79$\pm$0.02
11.41$\pm$0.04
11.36$\pm$0.02
11.35$\pm$0.04
14.43$\pm$0.03
2.34$\pm$0.06
0.23$\pm$0.02
3.46$\pm$0.04
13.33$\pm$0.11
12.12$\pm$0.03
14.23$\pm$0.04
Table 22: Stability results on the COMPAS dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
13.11$\pm$0.01
2.33$\pm$0.02
1.49$\pm$0.01
2.41$\pm$0.01
11.39$\pm$0.02
10.02$\pm$0.02
10.64$\pm$0.02
14.04$\pm$0.02
1.86$\pm$0.15
0.06$\pm$0.02
4.37$\pm$0.05
13.72$\pm$0.04
12.00$\pm$0.03
11.87$\pm$0.04
Table 23: Stability results on the GMSC dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
11.47$\pm$0.01
2.77$\pm$0.02
1.87$\pm$0.01
2.84$\pm$0.02
10.52$\pm$0.03
9.25$\pm$0.02
10.44$\pm$0.04
14.97$\pm$0.01
2.86$\pm$0.04
1.27$\pm$0.02
4.79$\pm$0.04
11.97$\pm$0.01
13.18$\pm$0.02
13.48$\pm$0.04
D.7 Remaining Results on ANN models
Table 24: Predictive faithfulness results on the Synthetic dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
@@ D.7 Remaining Results on ANN models
0.143$\pm$0.005
0.107$\pm$0.004
0.116$\pm$0.004
0.113$\pm$0.004
0.106$\pm$0.004
0.124$\pm$0.005
0.105$\pm$0.004
0.053$\pm$0.003
0.116$\pm$0.004
0.102$\pm$0.004
0.109$\pm$0.004
0.116$\pm$0.004
0.092$\pm$0.004
0.117$\pm$0.004
Table 25: Predictive faithfulness results on the German Credit dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.120$\pm$0.007
0.104$\pm$0.007
0.105$\pm$0.007
0.103$\pm$0.007
0.103$\pm$0.007
0.103$\pm$0.007
0.104$\pm$0.007
0.030$\pm$0.002
0.056$\pm$0.004
0.053$\pm$0.004
0.065$\pm$0.004
0.057$\pm$0.004
0.063$\pm$0.004
0.056$\pm$0.004
Table 26: Predictive faithfulness results on the HELOC dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.107$\pm$0.003
0.082$\pm$0.003
0.084$\pm$0.003
0.106$\pm$0.003
0.082$\pm$0.003
0.106$\pm$0.003
0.082$\pm$0.003
0.042$\pm$0.002
0.081$\pm$0.003
0.079$\pm$0.003
0.046$\pm$0.002
0.082$\pm$0.003
0.044$\pm$0.002
0.081$\pm$0.003
Table 27: Predictive faithfulness results on the Adult Income dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.207$\pm$0.004
0.071$\pm$0.003
0.070$\pm$0.003
0.221$\pm$0.004
0.070$\pm$0.003
0.223$\pm$0.004
0.071$\pm$0.003
0.069$\pm$0.003
0.230$\pm$0.004
0.233$\pm$0.004
0.071$\pm$0.003
0.230$\pm$0.004
0.068$\pm$0.003
0.230$\pm$0.004
Table 28: Predictive faithfulness results on the COMPAS dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.095$\pm$0.004
0.056$\pm$0.003
0.056$\pm$0.003
0.073$\pm$0.003
0.056$\pm$0.003
0.073$\pm$0.003
0.057$\pm$0.003
0.035$\pm$0.002
0.090$\pm$0.004
0.090$\pm$0.004
0.068$\pm$0.003
0.090$\pm$0.004
0.072$\pm$0.003
0.090$\pm$0.004
Table 29: Predictive faithfulness results on the GMSC dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.088$\pm$0.003
0.015$\pm$0.001
0.015$\pm$0.001
0.059$\pm$0.002
0.015$\pm$0.001
0.067$\pm$0.002
0.015$\pm$0.001
0.026$\pm$0.002
0.102$\pm$0.003
0.102$\pm$0.003
0.063$\pm$0.003
0.102$\pm$0.003
0.053$\pm$0.003
0.101$\pm$0.003
Table 30: Stability results on the Synthetic dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.58$\pm$0.00
5.29$\pm$0.20
3.56$\pm$0.17
5.21$\pm$0.18
31.27$\pm$1.00
10.55$\pm$0.02
21.11$\pm$0.89
10.80$\pm$0.01
4.23$\pm$0.22
2.51$\pm$0.20
4.13$\pm$0.20
30.62$\pm$1.00
9.74$\pm$0.02
20.60$\pm$0.91
15.12$\pm$0.04
5.47$\pm$0.48
5.79$\pm$0.39
5.33$\pm$0.37
32.06$\pm$1.00
13.94$\pm$0.04
21.70$\pm$0.82
Table 31: Stability results on the German Credit dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
12.84$\pm$0.00
3.18$\pm$0.31
2.16$\pm$0.33
3.19$\pm$0.31
11.57$\pm$0.06
11.80$\pm$0.01
11.77$\pm$0.05
12.77$\pm$0.01
3.06$\pm$0.36
1.98$\pm$0.34
3.06$\pm$0.36
11.47$\pm$0.06
11.77$\pm$0.02
11.73$\pm$0.05
16.06$\pm$0.06
3.24$\pm$0.43
4.34$\pm$0.36
3.96$\pm$0.21
14.40$\pm$0.11
15.13$\pm$0.07
14.78$\pm$0.10
Table 32: Stability results on the HELOC dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.74$\pm$0.00
3.53$\pm$0.17
2.47$\pm$0.07
3.69$\pm$0.23
10.96$\pm$0.08
10.69$\pm$0.02
10.42$\pm$0.05
11.52$\pm$0.01
3.00$\pm$0.23
2.03$\pm$0.09
3.18$\pm$0.29
10.85$\pm$0.07
10.54$\pm$0.02
10.32$\pm$0.05
15.16$\pm$0.04
4.49$\pm$0.50
4.10$\pm$0.18
5.32$\pm$0.33
13.91$\pm$0.06
14.32$\pm$0.04
14.02$\pm$0.08
Table 33: Stability results on the Adult Income dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
12.95$\pm$0.00
5.73$\pm$0.07
3.81$\pm$0.11
5.79$\pm$0.10
13.33$\pm$0.31
13.31$\pm$0.09
10.83$\pm$0.01
12.03$\pm$0.01
3.98$\pm$0.16
1.97$\pm$0.13
4.02$\pm$0.15
13.98$\pm$0.39
12.01$\pm$0.09
9.67$\pm$0.02
14.75$\pm$0.05
3.58$\pm$0.33
3.38$\pm$0.07
4.20$\pm$0.17
17.08$\pm$0.45
14.49$\pm$0.12
12.95$\pm$0.06
Table 34: Stability results on the COMPAS dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
13.11$\pm$0.01
4.68$\pm$0.18
3.59$\pm$0.04
4.99$\pm$0.30
17.33$\pm$0.13
10.55$\pm$0.02
15.23$\pm$0.10
12.88$\pm$0.01
3.96$\pm$0.19
2.83$\pm$0.05
4.30$\pm$0.32
17.21$\pm$0.12
10.39$\pm$0.01
15.13$\pm$0.09
15.30$\pm$0.03
4.30$\pm$0.53
3.86$\pm$0.30
5.46$\pm$0.42
18.09$\pm$0.10
12.80$\pm$0.04
16.37$\pm$0.11
Table 35: Stability results on the GMSC dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.47$\pm$0.01
4.44$\pm$0.09
3.63$\pm$0.02
4.59$\pm$0.11
10.35$\pm$0.03
9.94$\pm$0.02
9.77$\pm$0.03
11.51$\pm$0.01
3.78$\pm$0.11
2.94$\pm$0.03
3.96$\pm$0.14
10.28$\pm$0.03
9.94$\pm$0.01
9.76$\pm$0.03
14.72$\pm$0.04
4.22$\pm$0.55
4.40$\pm$0.26
5.57$\pm$0.48
13.29$\pm$0.05
12.73$\pm$0.06
13.65$\pm$0.05
[Uncaptioned image] Figure 3: Fairness analysis of PGU metric on the German Credit dataset with ANN model. Shown are average and standard error values for majority (male) and minority (female) subgroups. Larger gaps between the values of majority and minority subgroups (i.e, red and blue bars respectively) indicate higher disparities which are undesirable. [Uncaptioned image] Figure 4: Fairness analysis of PGU metric on the Adult Income dataset with ANN model. Shown are average and standard error values for majority (male) and minority (female) subgroups. Larger gaps between the values of majority and minority subgroups (i.e, red and blue bars respectively) indicate higher disparities which are undesirable.
Appendix E Choice of XAI methods, datasets, and models
While feature attribution-based explanation methods such as LIME, SHAP, and Gradient-based methods have been proposed a few years back, they continue to be the most popular and widely used post hoc explanation methods both in research [Chen et al., 2022, Krishna et al., 2022, Alvarez-Melis and Jaakkola, 2018, Hooker et al., 2018, Han et al., 2022] and in practice [Elshawi et al., 2019, Whitmore et al., 2016, Ibrahim et al., 2019, Ghassemi et al., 2021]. In fact, several recent works published in 2022 have analyzed these methods both theoretically and empirically, and have called for further study of these methods given their widespread adoption [Dasgupta et al., 2022, Chen et al., 2022, Balagopalan et al., 2022, Han et al., 2022, Krishna et al., 2022, Fokkema et al., 2022]. Furthermore, recent research has also argued that there is little to no understanding of the behavior and effectiveness of even basic post hoc explanation methods such as LIME, SHAP, and gradient-based methods [Krishna et al., 2022, Lipton, 2016, Rudin, 2019, Kaur et al., 2020, Bansal et al., 2020, Agarwal and Nguyen, 2020], and developing such an understanding would be a critical first step towards the progress of the XAI field. To this end, we focus on these methods for the first release of our OpenXAI framework. In the next release, we plan to evaluate and benchmark other recently proposed methods (e.g., TCAV and its extensions, influence functions, etc.) as well.
Note that the evaluation metrics and the explanation methods that we include in our framework are generic enough to be applicable to other modalities of data including text and images. The reason why we focused on tabular data for the first release of OpenXAI is two-fold: i) The need for model understanding and explainability is often motivated by high-stakes decision-making settings and applications e.g., loan approvals, disease diagnosis and treatment recommendations, recidivism prediction etc. [Borisov et al., 2021, Pawelczyk et al., 2020, 2021]. Data encountered in these settings is predominantly tabular. ii) Recent research has argued that there is no clear understanding as to which explanation methods perform well on what kinds of metrics even on simple, low-dimensional tabular datasets [Lipton, 2016, Dasgupta et al., 2022, Krishna et al., 2022, Han et al., 2022], and that this is a big open question which has far reaching implications for the progress of the field. To this end, we focused on tabular data for the first release of OpenXAI so that we could find some answers to the aforementioned question in the context of simple, low-dimensional tabular datasets before proceeding to high-dimensional image and text datasets. In the next release of OpenXAI, we plan to include and support image and text datasets, and also add metrics and explanation methods that are specific to these new data modalities.
The datasets that we utilize in this work are very popular and are widely employed in XAI and fairness research till date. For instance, several recent works in XAI published at ICML, NeurIPS, and FAccT conferences in 2021-22 have employed these datasets both to evaluate the efficacy of newly proposed methods, as well as to study the behavior of existing methods [Balagopalan et al., 2022, Dai et al., 2022, Dasgupta et al., 2022, Dominguez-Olmedo et al., 2022, Karimi et al., 2020a, Slack et al., 2021, Upadhyay et al., 2021, Agarwal et al., 2021]. Given this, we follow suit and employ these datasets in our benchmarking efforts. Similarly, the aforementioned works also employ logistic regression models and deep neural network architectures similar to the ones considered in our research.